NAOJ GW Elog Logbook 3.2

Michael and Yuhang
When IR phase shifter (IRPS) is driven, an obvious misalignment could be found in the IRMC scanning spectrum. In this elog, we report the detail of misalignment when IRPS is given different level of high voltage.
Experimental setup: IRMC is scanned to have two TEM00 peaks within one slope of ramp signal. The IRPS is located before IRMC and given an 75V high voltage at the beginning. We changed the high voltage level and took scanning spectrum accordingly. The measurements were done with the following high voltage
High voltage |
75V |
75V (zoom in) |
65V (zoom in) |
55V (zoom in) |
45V (zoom in) |
35V (zoom in) |
25V (zoom in) |
15V (zoom in) |
5V (zoom in) |
We got scanning spectrum as attached figure 1. It is clear from figure 1 that almost only TEM01 mode appears after IRPS is given different high voltage.
According to the peak values in figure 1, we extract the percentage of power goes to TEM01 as figure 2.

Aritomi, Marc
Following the discovery of the not fixed lens on the IR path, we started to realign this beam.
First, I fixed by hand this lens at a position where the AC signal with the surface reference sample was ~0.08 (to compare to the previous ~0.25 and expected 0.36).
Then, we made sure the beam was horizontal after the 2 lenses using the pinhole at 45 and 27 mm on Z.
This allowed us to measure the beam power transmitted by the pinhole at various Z position.
From this measurement, it was possible to extract the waist size (~70um ) and position (~34 mm).
I took 2 points for the red beam to find the crossing point of these 2 beams.
And it can be seen (fig 1) that the ir beam has to be shifted horizontally by 232 um to have the crossing point at the the ir waist.
I started to move the ir beam using again the pinhole at 2 positions but it was required to move this beam both horizontally and vertically.
I'll finish this alignment this morning.
Note that compared to the previous situation, the crossing point is 3 mm closer towards the periscope.

Yuhang and Michael
We refer to the previous logbook entries:
1904 - Reference measurement to IRPS jitter noise
2393 - First entry on this topic describing the new layout of the phase shifter to orient in perpendicular incidence and jitter noise
2407 - Previous entry on IRPS replacement, with measurement of X and Y channels of PSD
This time, we performed a series of measurements of the X and Y channels on the PSD, divided by the T (total) channel (i.e. frequency response 2/1 in spectrum analyser). This was motivated by the entries 1904 and 2407, where after discussion we determined it was ambiguous whether or not we were measuring jitter or amplitude noise - since we were measuring past the IRMC, we are taking the transmission of the mode cleaner, the power output of which is affected by the alignment of the input beam. This also motivated Yuhang to simulate the effect of beam waist positioning on the noise at the PSD past the IRMC.
First, we took the measurements of X/T and Y/T using a PSD just past the homodyne detector's flipping mirror, shown in figure 1 and 2. In the X/T (Yaw) case, the relative contribution of the X and T channels changes very little with the amount of excitation. The noise floor is similar to no excitation with the difference of a broad peak at about 2.3 kHz, which could correspond to the beam vibration frequency of the PZT element supporting the phase shifter mirror. In the Y/T case (pitch), the contribution of the Y channel actually decreases with respect to T. However, we saw in 2407 that the absolute value of the Y noise is quite high. This indicates that there is quite a lot of amplitude noise introduced on the Y channel. Yuhang's simulation indicates that this amplitude noise may be caused by the phase shifter being offset from the beam waist. The results motivate us to do the following two measurements, prior to adjusting the relative position of waist/phase shifter: 1 - measure the X, Y/T noise induced by the phase shifter excitation, with the PSD before the IRMC, and 2 - measure the spectrum of the IRMC when changing the voltage sent from its high voltage driver, specifically looking at the behaviour of higher order modes as the IRMC is mismatched.
So far, we have taken measurements of the jittering before the IRMC. A sketch is shown in figure 3. Measurements are shown in figures 4-6 (some traces at certain values of excitation are missing due to data corruption) In this layout, we are definitely measuring more angular deflection now. In all cases,the contribution of X and Y increases relative to T with increased phase shifter excitation, and is also higher in relative magnitude, often above 0 dB. By contrast, the X, Y/T do not go above 0 dB after the IRMC. This increased noise may also be due to some other broad resonances at about 12 kHz and 30 kHz.

Yuhang and Michael
We will assemble the new OPO in the ATC cleanroom. We wanted to see if the laser beam set up in the corner is collimated. We measured the beam using the beam profiler from TAMA FDS cleanroom, with z = 0 being the position of the last lens that was fixed on the bench as we got here (figure 1).
Using curve fitting, we find the following fit of the beam size (figure 2), where blue refers to the horizontal axis and orange the vertical axis. The beam is not collimated and also a bit astigmatic. The fitting on the horizontal size also is quite distant from the data point closer to z = 0.
Comment on the characterisation of ATC cleanroom laser.
The beam profiler output shown in attached figures. We can see the beam position on the profiler for each measurement. The number on Gaussian -> 13.5% is taken as the beam diameter. V is the vertical axis and W is the horizontal axis.
Axis of figure 2 should be "beam radius"

Yesterday I acted on the 2 lenses on the ir path to shift the beam vertically and superpose it to the red.
I placed the pinhole on the translation stage and moved it back and forth along the optical axis to maximize the power on the power-meter.
Then I did a scan of the surface reference sample that showed little improvement (R~12.5).
Matteo noticed that the crossing of the 2 beams did not correspond to the ir waist position.
I wanted to move the lenses on the ir path lateraly to shift the beam.
However, I found out that the first lens was not fixed at all : the mount got easily out of the fork and I had to unscrew the fork to fix the lens...
I tried my luck to recover a good position of this lens by hand but it is too sensitive (even though for a brief instant I could see R~18 which is the expected value, it was not possible to fix it at this position alone) so I'll have to restart the ir alignment...

Marc, Michael, Yoichi, Yuhang
Due to the earthquake reported in elog2416, we had issue of PR/BS pico-motors. To fix this problem, we opened PR/BS chambers today.
In the end, we fixed problems of PR/BS position and pico-motors. PR/BS chambers have been closed. But the air was not evacuated. Probably, we can evacuate on tomorrow.
While PR/BS chambers are open, we found issues as following:
1. PR pitch pico-motor is close to the end of range. (Figure 1)
2. BS mirror is too low, which makes the upper horizontal earthquake stop not useable.
3. BS earthquake stop is too far from mirror. The distance was reduced to be around 1mm now.

We can send infrared light from input-coupler side. After that, we take reflection and transmission power. By doing ring-down measurement, we can extrapolate well information of P0, P1, r1, t1.
P0 is power coupled to OPO cavity, P1 is power not-coupled to OPO cavity, r1 is amplitude reflectivity of input coupler, t1 is amplitude transmissivity of input copuler.
I did a calculation with the ideal parameters of OPO, which tells us decay time of ~4us.
To measure this decay, we should be able to lock OPO and switch off incident laser faster than ~100ns. To do this we need to use signal generator to send a square wave to an AOM which is before OPO. According to the spec of MT110-A1.5-1064, the rise time can be smaller than 100ns if the beam size is smaller than 0.6mm (diameter). We will design a small enough beam to make this rise time small enough to measure ring-down.
The expected ring-down for reflection and transmission is attached as figure 1.
Since OPO is finally closed, the next step is to characterize the intra-cavity losses. This is important for us because we are suspecting some of the optical losses are from OPO (current estimated loss budget for FDS). So this is an important step to understand the loss budget in the frequency dependent squeezing experiment.
I modified a bit the code to see the difference of measurement for different OPO intra-cavity losses.
Now, the laser is injected from the crystal side of OPO. I did a simulation of this case as Fig. 1. In this case, we will miss the information of OPO reflection. The blue and orange curves overlap for reflection.
If laser is injected from in-coupling mirror, as shown in Fig.2, we find that although decay time is not enough to indicate optical losses. We can extract losses from reflection signal.
So we will rotate OPO next week and inject laser from the in-coupling mirror side.

Using measurements of entry 2391 I could compare the IR and red beams propagation directions as in the attached figure (axis unit in mm):
- the angle the 2 beam : ~0.07rad
- the height difference between the 2 beams : ~ 100 um (might explain the difficulties to align)
I'll try to use the pinhole to shift the IR beam height to the red beam one.

Earth quake happened on 20210320 18:09, which was quite strong and long.
We checked the oplev signal of all suspended mirrors (attached figure 2). The watchdog of BS and END mirrors were switched off. However, since we didn't give large offset, we can just use the oplev signal to check how much mirror drifted. We could see these change
Yaw_change | Pitch_change | |
PR | -100 | +2500 |
BS | -500 | +1040 |
Input | -320 | +2450 |
End | +720 | -4100 |

Pierre and Yuhang
To investigate if SNR can be improved by an improved photo detector design, we conducted TAMA resonant PD simulation by using two simulation tools. One is a python code called 'zeros', the other is a Texas Instrument software 'TINA'.
The configuration is based on the scheme in this link. In the real case, we have modified this scheme, which results in two cases. One case is using the same op-amp with R1 10Ohm and R2 100Ohm. The other case is using op-amp LMH6624 with R1 1kOhm and R2 10kOhm.
To measure PD noise, we used a 32dB amplifier to make sure the PD noise is well above the instrument noise. The use of this amplifier has been considered also in the simulation software. In the first attached figure, there are electronic noise measurement and simulation results. We could see that python code simulates well the floor noise. However, the noise around 14MHz is better simulated by TINA.
We anticipated a signal of 0.11uA from CCFC field. To compare signal and noise, we divided the voltage noise from figure one by the PD gain. Then we get the current input noise as the attached figure 2. This DC value of signal is much higher than the noise level at 14MHz. From the python simulation, LMH6624 has better performance. However, simulation of TINA tells us that AD8057 has better performance.
To compare, we put here also a measurement of PD noise budget (figure 3). From this noise budget, TINA simulation is closer with the real measurement.
Figure 4 and 5 show the SNR of different PD configurations with TINA simulation

Marc, Yuhang
Following the End mirror OpLev check (entry 2412) and the measurement with the sensing matrix of entry 2411 we decided to use a 4x4 matrix to try to remove the coupling of unwanted degrees of freedom on the reconstructed signals from the 2 QPDs.
- We locked the FC and only close End mirror length control.
- Tune the phases : We found out that all the phases changed by ~40 degrees. So we used time series of all segments and maximized the I signals
-
We sent a line at 2 Hz on Input and End pitch and yaw. Using the calibration of entry 1877 we injected a 3.78 mrad line for every dof. Namely
Pitch Yaw Input 600 172 End 727 141 - We measured the amplitude of the injected line (removing the background level noise) on each QPDs.
- We computed the driving matrix and had to modify the medm configuration (outmatrix.adl) to allow to use the 4x4 matrix
-
QPD1 pitch QPD2 pitch QPD1 yaw QPD2 yaw Input pitch 6 -5 -0.5 -1 End pitch -5 10 0.5 1 Input yaw 1 -2 4 -3 End yaw 2 -4 -4 12 - We checked each QPDs signals as in entry 2412 but coupling was still visible.
- We checked the demodulation phases and they seem to have moved by ~-40 degrees back to the previous phases...
- We checked possible reasons for this phase changes :
- We aligned the FC to another position : no phase changes
- We tried to put the optical bench to the ground (before putting the electronics inside of a clean room the optical table was connected to the nim racks ground) : no changes
- We putted back the 2x2 matrix of entry 2412 and tried to tuned the coupled coefficients by hand to remove coupling. It could approximatively work but during this tuning we could also see that the level of coupling increases or decreases sometimes without noticeable reasons...

Marc, Yuhang
We went to check to End mirror OpLev : All optics seemed well fixed.
We tuned the beam height to be identical on the injection and detection windows.
We moved vertically the lens on the detection window to have the beam centered.
Unfortunately it didn't change the coupling of pitch to yaw visible in the figure 1 of the parent entry.
We can see that pitch couples to yaw but not yaw to pitch.
This seems to indicate that the sensing matrix of OpLev is not the culprit.
Furthermore, when we inject a line on the End mirror pitch, we can indeed see the beam on the FC transmission camera moving in pitch but not in yaw. Even if the OpLev yaw senses this line...
On input mirror we can see both pitch and yaw lines on pitch and yaw signals from the oplev so it could be solved by finer tuning of the oplev rotation matrix.

Marc and Yuhang
We sent several sine wave to different DOF for filter cavity as following:
2Hz | 3Hz | 9Hz | 5Hz |
Input pitch | Input Yaw | End Pitch | End Yaw |
(At the beginning, 2Hz/3Hz/5Hz/7Hz were used. We will use 7Hz instead of 9Hz in the future.)
By sending these sine waves to different DOF, the check of coupling in reconstruction becomes easier. As shown in the attached figure, we checked many signals before and after excitation. They are:
Oplev signals: Red line: Oplev signal before excitation Green line: Oplev signal after excitation
AA signals: Blue line: AA noise when light goes to QPDs are blocked (actually this is DGS ADC noise) Brown line: AA reconstructed signal before excitation Pink line: AA reconstructed signal after excitation
(AA signal is not calibrated, so it is much larger than oplev signal)
Oplev issues:
Since a large 9Hz peak appears in the oplev signal of End mirror yaw, either End mirror pitch driving or End mirror yaw sensing is strange.
AA issues:
End mirror pitch signal couples to End and Input Yaw.
Input mirror pitch signal couples to End Yaw.
Plan:
1. We are going to check end mirror oplev and coils. 2. Consider to realize a matrix including pitch/yaw coupling
Marc, Yuhang
We went to check to End mirror OpLev : All optics seemed well fixed.
We tuned the beam height to be identical on the injection and detection windows.
We moved vertically the lens on the detection window to have the beam centered.
Unfortunately it didn't change the coupling of pitch to yaw visible in the figure 1 of the parent entry.
We can see that pitch couples to yaw but not yaw to pitch.
This seems to indicate that the sensing matrix of OpLev is not the culprit.
Furthermore, when we inject a line on the End mirror pitch, we can indeed see the beam on the FC transmission camera moving in pitch but not in yaw. Even if the OpLev yaw senses this line...
On input mirror we can see both pitch and yaw lines on pitch and yaw signals from the oplev so it could be solved by finer tuning of the oplev rotation matrix.

Marc and Yuhang
We have checked the important parameters and green power as following:
SHG temperature: 3.09kOhm
AOM setting: 109.036MHz, 5.5dBm
green power before AOM: 50.1mW
green power before FC injection: 24.9mW

Yuhang and Michael
Unfortunately in 2398 the measurement was performced incorrectly. Random noise from the spectrum analyser was injected into the "Perturb IN" port of the IRMC servo controller rather than the IN port of the IRMC HVD. This mistake was corrected, and we measured the following
- Power spectrum of dark noise, LO jitter and phase shifter excitation noise from the X, Y and T (Total) channels of the 500 Ohm position sensitive detector (PSD) - figures 1, 2 and 3 respectively
- Transfer functions and coherence of source noise to the X, Y and T channels of the PSD - figure 4
- Transfer functions of source noise to the IRMC error signal (at EPS1), correction signal (at Servo Out) and IRMC reflection - figure 5
The high frequency numbers for pitch are now much more in line with 1904. Notably, the mid frequency discrepancy between 2398 and 1904 still remains, where the mid frequency noise magnitude in this rearranged configuration is much higher than for 1904 despite having similar levels at the extremes of the measurement window. The PSD of yaw noise from the excited phase shifter is now signifcantly reduced compared to 1904, although the LO jitter is slightly above the PSD dark noise this time. We also measured the T-channel of the PSD to obtain the amplitude noise of the LO beam.
There is quite a lot of coherence from the source noise to both T-channel noise and pitch noise, especially amplitude noise above 30 Hz. The T-channel noise transfer function and coherence is fairly similar to the IR reflection, which is to be expected since they essentially just amplitude noise transmitted and reflected from the IRMC.
Source noise coupling to the error signal increases above 100 Hz, while the correction signal transfer function does not follow said increase. Howver, they have similar coherence across all frequencies.
Just as a note, when I checked and realigned the IRMC at the start of the measurements for the day, it seemed to have drifted out of alignment with a pitch HOM prominent in the reflection spectrum.

Marc and Yuhang
Marc provided us a code written by Julia Casanueva Diaz, which was used by Marc to do FIS auto alignment for Virgo.
After reading this code together, I found out what can be improved for the construction of driving matrix. Especially, I finally understand what is the meaning of matrix of inversion. The idea is the change of basis after inversion. With this in mind, we checked again the sensing matrix and inverted it. We got following driving matrix
|
Pitch |
Yaw |
||
|
WFS1 |
WFS2 |
WFS1 |
WFS2 |
Input mirror |
7.37 |
-4.98 |
-1.71 |
1.65 |
End mirror |
6.64 |
9.69 |
-1.65 |
-4.41 |
With this driving matrix, we did preliminary check of Input and End mirror pitch/yaw reconstructed spectrum. Compared with the old measurement (elog2263, elog2245), it seems the reconstruction is more reasonable (attached figure).
More investigation and optimization will be done soon.

Marc and Yuhang
After moving picomotors, filter cavity alignment was recovered.
Spectrum of all mirrors were checked as the attached figure. No touching induced peaks were found.

For the Labview control for INPUT mirror picomotor, I have some comments.
***notice: check point of reflected beam is inside PR chamber
1. Pitch positive is to move the reflected beam to the left.
2. Yaw positive is to move the relfected beam to the up.
(currently, the input mirror pitch and yaw picomotors are swapped)

As reported in entry 2399, there was an earthquake in March 16th around 5 am.
I went to check the setup and it seems that at least the imaging unit is misaligned (moved by few 10 um in distance to the translation stage and the red beam was misaligned by ~ 1 cm to the left on the photodiode). I'll try to characterize the beam again after maximizing the red and IR beams on the photodiode/powermeter.
Also, I suspect that there was some troubles with the computer as :
- several errors messages on the LabView program referring to troubles with the translation stage
- waveplate moved around the earthquake time but no errors
- the chopper frequency changed to 1kHz.

Michael and Yuhang
We completed the setup as per 2393. The spectrum of aligment noise from excitation of the IRPS was made and compared to 1904.
The measurement was taken using a 500 Ohm PSD placed after the IRMC and homodyne flipping mirror. We measure the dark noise, followed by the jitter noise when transmitting the local oscillator, and then the jitter noise when the phase shifter PZT is excited by random noise from the spectrum analyser, applied via the high voltage driver. In the original measurement of 1904, the PZT was excited using 200 mVpk random noise, where the beam had a horizontal angle of incidence of ~ 45 degrees. Here, the beam has perpendicular incidence, so we only use 200/sqrt(2) ~ 140 mVpk excitement for the yaw measurement, but keep 200 mVpk for the pitch measurement.
For yaw, we seen that the noise floor seems to be the same or perhaps slightly better than 1904, it is hard to tell just from the picture. However, the new measurement reaches the noise floor at approximately 30 Hz, compared to ~ 10 Hz in the old measurement. Also, both measurements have similar values at either end of the measurement frequency window, but the new measurement looks to be a lot higher in the mid frequencies than that of the old measurement. The PD dark noise is consistent in both cases.
For pitch, the measurement is very strange. Unlike yaw in this case and both degrees in 1904, the LO jitter noise does not converge to the PSD dark noise at high frequency. Both the LO jitter and excitation jitter seem to be about 15-20 dB higher than their yaw counterparts.
Afterwards, we measured the transfer function of input excitation (from the spectrum analyser) to measured PSD noise. We see that the pitch transfer function has significantly more coherence from the excitation to the PSD. The yaw coupling at 140 mVpk excitation is also much lower than for pitch excitations.
A few other notes:
- The IRMC alignment seems to drift quite a lot. Last week it had to be relocked about 4 or 5 times during the afternoon. It also had dropped to about 1.6 mW transmitted power before starting the measurement described in this elog. I realigned it for the optimal transmission each time (1.8 mW).
- The phase shifter is not quite flat, it has some pitch tilt. As a result, the incoming alignment is not perfectly horizontal, but rather is adjusted so that the outgoing beam maintains a constant height of 76 mm.
- PSD output voltage was zeroed to within 0.5 mV for both X and Y